Computational Algebraic Geometry Lab

Brendan Hassett, Amanda Knecht, Ning Song

We will introduce algorithms and computer packages commonly used in algebraic geometry. These can be applied to obtain computer verifications of some theorems from high school geometry, culminating in the Descartes circle theorem. While brute-force computations are not always the most elegant proofs, they are a powerful tool for evaluating research hypotheses in algebraic geometry.